# Holographic thermodynamic relation for dissipative and non-dissipative universes in a flat FLRW cosmology

### v1

@@ I Introduction
To explain the accelerated expansion of the late Universe PERL1998_Riess1998 ; Planck2018 ; Hubble2017 , astrophysicists have proposed various cosmological models, e.g., lambda cold dark matter ($\Lambda$CDM) models, time-varying $\Lambda(t)$ cosmology FreeseOverduin ; Nojiri2006etc ; Valent2015Sola2019 ; Sola_2009-2022 , creation of CDM (CCDM) models Prigogine_1988-1989 ; Lima1992-1996 ; LimaOthers2023 ; Freaza2002Cardenas2020 , bulk viscous cosmology BarrowLima ; BrevikNojiri ; EPJC2022 , and thermodynamic scenarios such as entropic cosmology EassonCai ; Basilakos1 ; Koma45 ; Koma6 ; Koma7 ; Koma8 ; Koma9 ; Neto2022 ; Gohar2024 . These studies imply that our Universe should finally approach a de Sitter universe whose horizon is considered to be in thermal equilibrium. The thermodynamics of the universe has been examined from various perspectives, e.g., the first law of thermodynamics Cai2005 ; Cai2011 ; Dynamical-T-2007 ; Dynamical-T-20092014 ; Sheykhi1 ; Sheykhi2Karami ; Santos2022 ; Sheykhia2018 ; ApparentHorizon2022 ; Cai2007 ; Cai2007B ; Cai2008 ; Sanchez2023 ; Nojiri2024 ; Odintsov2023ab ; Odintsov2024 ; Mohammadi2023 ; Odintsov2024B , the second law of thermodynamics Easther1-Egan1 ; Pavon2013Mimoso2013 ; Bamba2018Pavon2019 ; deSitter_entropy ; Saridakis20192021 , and the holographic equipartition law related to the emergence of cosmic space Padma2010 ; Verlinde1 ; HDE ; Padmanabhan2004 ; ShuGong2011 ; Koma14 ; Koma15 ; Koma16 ; Koma17 ; Koma19 ; Koma20 ; Padma2012AB ; Cai2012 ; Hashemi ; Moradpour ; Wang ; Koma10 ; Koma11 ; Koma12 ; Koma18 ; Krishna20172019 ; Mathew2022 ; Chen2022 ; Luciano ; Mathew2023 ; Mathew2023b ; Pad2017 ; Tu2018 ; Tu2019 .
@@ II.2 Entropy $S_{H}$ and temperature $T_{H}$ on the horizon
The second law of thermodynamics on the horizon, $\dot{S}_{\rm{BH}}\geq 0$, is satisfied in favored cosmological models, such as $\Lambda$CDM models Koma14 . In the present study, the form of the Bekenstein–Hawking entropy, $S_{\rm{BH}}$, is typically used for the entropy $S_{H}$ on the cosmological horizon. (Various forms of black-hole entropy such as nonextensive entropy have been proposed Das2008 ; Radicella2010 ; LQG2004_123 ; Tsallis2012 ; Czinner1Czinner2 ; Barrow2020 ; Nojiri2022 , as described in Refs. Koma18 ; Koma19 ; Koma20 . For a general form of entropy related to the Friedmann equation, see, e.g., Ref. Nojiri2024 .)
@@ III First law of thermodynamics and the modified thermodynamic relation
In this study, we typically use $S_{H}=S_{\rm{BH}}$ and $T_{H}=T_{\rm{KH}}$. However, $S_{H}$ and $T_{H}$ are retained so that we can consider other entropies and temperatures. For example, when nonextensive entropies Das2008 ; Radicella2010 ; LQG2004_123 ; Tsallis2012 ; Czinner1Czinner2 ; Barrow2020 ; Nojiri2022 are used for $S_{H}$, the first law of thermodynamics should lead to the modified Friedmann and acceleration equations, as examined in previous works; see, e.g., Ref. Odintsov2024 . Also, a general form of entropy has been discussed based on the first law Nojiri2024 .
@@
The additional $\dot{f}_{\Lambda}(t)$ and $h_{\textrm{B}}(t)$ terms in Eq. (29) are based on two extra driving terms included in general cosmological equations, and these two terms are assumed to be related to entropies. In this paper, the $\dot{f}_{\Lambda}(t)$ term is considered to be related to the reversible entropy $S_{\dot{f}_{\Lambda}}$, such as that related to the reversible exchange of energy. The $h_{\textrm{B}}(t)$ term is considered to be related to the irreversible entropy $S_{h_{\textrm{B}}}$ arising from dissipative processes, such as the bulk viscosity and the creation of CDM Koma7 . Other entropies, such as nonequilibrium entropies and generalized entropies, are not considered here. The relationship between similar additional terms and nonequilibrium entropies has been discussed in, e.g., Refs. Dynamical-T-2007 ; Cai2007B ; Santos2022 . A general form of entropy that connects the Friedmann equations for any gravity theory with horizon thermodynamics has been examined in Ref. Nojiri2024 . We expect that modified FLRW equations can be derived from Eq. (29), using various forms of entropy $S_{H}$ on the horizon Das2008 ; Radicella2010 ; LQG2004_123 ; Tsallis2012 ; Czinner1Czinner2 ; Barrow2020 ; Nojiri2022 . The modified thermodynamic relation given by Eq. (29) should provide new insights into thermodynamics and cosmological equations.
@@ IV.2 Applying the equipartition law of energy on the horizon to the modified thermodynamics relation
In fact, the holographic entanglement entropy RyuTakayanagi2006 is equivalent to the formula for the entropy on a cosmological horizon in a de Sitter space Arias2020 . In addition, gravity should be related to the relative entropy corresponding to the free-energy difference Relative_entropy , based on the holographic entanglement entropy. Accordingly, the free-energy difference $dF_{H}$ included in the modified thermodynamic relation may be related to gravity through the relative entropy and the holographic entanglement entropy. These tasks are left for future research.
@@ V Evolution of thermodynamic quantities in a constant $T_{\rm{KH}}$ model
In this section, we observed typical evolutions of the thermodynamic quantities in the modified thermodynamic relation using the $\Lambda$CDM model and the constant $T_{\rm{KH}}$ model. The modified thermodynamic relation should help to understand the properties of various cosmological models from a thermodynamic viewpoint. The present study should provide new insights into the discussion of thermodynamic cosmological scenarios.

### v2

@@ I Introduction
To explain the accelerated expansion of the late Universe PERL1998_Riess1998 ; Planck2018 ; Hubble2017 , astrophysicists have proposed various cosmological models, e.g., lambda cold dark matter ($\Lambda$CDM) models, time-varying $\Lambda(t)$ cosmology FreeseOverduin ; Nojiri2006etc ; Valent2015Sola2019 ; Sola_2009-2022 , creation of CDM (CCDM) models Prigogine_1988-1989 ; Lima1992-1996 ; LimaOthers2023 ; Freaza2002Cardenas2020 , bulk viscous cosmology BarrowLima ; BrevikNojiri ; EPJC2022 , and thermodynamic scenarios such as entropic cosmology EassonCai ; Basilakos1 ; Koma45 ; Koma6 ; Koma7 ; Koma8 ; Koma9 ; Neto2022 ; Gohar2024 . These studies imply that our Universe should finally approach a de Sitter universe whose horizon is considered to be in thermal equilibrium. The thermodynamics of the universe has been examined from various perspectives, e.g., the first law of thermodynamics Cai2005 ; Cai2011 ; Dynamical-T-2007 ; Dynamical-T-20092014 ; Sheykhi1 ; Sheykhi2Karami ; Santos2022 ; Sheykhia2018 ; ApparentHorizon2022 ; Cai2007 ; Cai2007B ; Cai2008 ; Sanchez2023 ; Nojiri2024 ; Odintsov2023ab ; Odintsov2024 ; Mohammadi2023 ; Odintsov2024B , the second law of thermodynamics Easther1-Egan1 ; Pavon2013Mimoso2013 ; Bamba2018Pavon2019 ; deSitter_entropy ; Saridakis2019 ; Saridakis2021 ; Sharif2024 , and the holographic equipartition law related to the emergence of cosmic space Padma2010 ; Verlinde1 ; HDE ; Padmanabhan2004 ; ShuGong2011 ; Koma14 ; Koma15 ; Koma16 ; Koma17 ; Koma19 ; Koma20 ; Padma2012AB ; Cai2012 ; Hashemi ; Moradpour ; Wang ; Koma10 ; Koma11 ; Koma12 ; Koma18 ; Krishna20172019 ; Mathew2022 ; Chen2022 ; Luciano ; Mathew2023 ; Mathew2023b ; Pad2017 ; Tu2018 ; Tu2019 ; Chen2024 .
@@ II.2 Entropy $S_{H}$ and temperature $T_{H}$ on the horizon
The second law of thermodynamics on the horizon, $\dot{S}_{\rm{BH}}\geq 0$, is satisfied in favored cosmological models, such as $\Lambda$CDM models Koma14 . In the present study, the form of the Bekenstein–Hawking entropy, $S_{\rm{BH}}$, is typically used for the entropy $S_{H}$ on the cosmological horizon. (Various forms of black-hole entropy such as nonextensive entropy have been proposed Das2008 ; Radicella2010 ; LQG2004_123 ; Tsallis2012 ; Czinner1Czinner2 ; Barrow2020 ; Nojiri2022 ; Gohar2023 , as described in Refs. Koma18 ; Koma19 ; Koma20 . For a general form of entropy related to the Friedmann equation, see, e.g., Ref. Nojiri2024 .)
@@ III First law of thermodynamics and the modified thermodynamic relation
In this study, we typically use $S_{H}=S_{\rm{BH}}$ and $T_{H}=T_{\rm{KH}}$. However, $S_{H}$ and $T_{H}$ are retained so that we can consider other entropies and temperatures. For example, when nonextensive entropies Das2008 ; Radicella2010 ; LQG2004_123 ; Tsallis2012 ; Czinner1Czinner2 ; Barrow2020 ; Nojiri2022 ; Gohar2023 are used for $S_{H}$, the first law of thermodynamics should lead to the modified Friedmann and acceleration equations, as examined in previous works; see, e.g., Ref. Odintsov2024 . Also, a general form of entropy has been discussed based on the first law Nojiri2024 .
@@
The additional $\dot{f}_{\Lambda}(t)$ and $h_{\textrm{B}}(t)$ terms in Eq. (29) are based on two extra driving terms included in general cosmological equations, and these two terms are assumed to be related to entropies. In this paper, the $\dot{f}_{\Lambda}(t)$ term is considered to be related to the reversible entropy $S_{\dot{f}_{\Lambda}}$, such as that related to the reversible exchange of energy. The $h_{\textrm{B}}(t)$ term is considered to be related to the irreversible entropy $S_{h_{\textrm{B}}}$ arising from dissipative processes, such as the bulk viscosity and the creation of CDM Koma7 . Other entropies, such as nonequilibrium entropies and generalized entropies, are not considered here. The relationship between similar additional terms and nonequilibrium entropies has been discussed in, e.g., Refs. Dynamical-T-2007 ; Cai2007B ; Santos2022 . A general form of entropy that connects the Friedmann equations for any gravity theory with horizon thermodynamics has been examined in Ref. Nojiri2024 . We expect that modified FLRW equations can be derived from Eq. (29), using various forms of entropy $S_{H}$ on the horizon Das2008 ; Radicella2010 ; LQG2004_123 ; Tsallis2012 ; Czinner1Czinner2 ; Barrow2020 ; Nojiri2022 ; Gohar2023 . The modified thermodynamic relation given by Eq. (29) should provide new insights into thermodynamics and cosmological equations.
@@ IV.2 Applying the equipartition law of energy on the horizon to the modified thermodynamics relation
In fact, the holographic entanglement entropy RyuTakayanagi2006 is equivalent to the formula for the entropy on a cosmological horizon in a de Sitter space Arias2020 . In addition, gravity should be related to the relative entropy corresponding to the free-energy difference Relative_entropy , based on the holographic entanglement entropy. The two entropies are usually discussed in a universe at constant temperature. Accordingly, the free-energy difference $dF_{H}$ included in the modified thermodynamic relation may be related to gravity through the relative entropy and the holographic entanglement entropy. These tasks are left for future research.
@@ V Evolution of thermodynamic quantities in a constant $T_{\rm{KH}}$ model
The constant $T_{\rm{KH}}$ model is reviewed, according to a previous work Koma19 . Systems at constant temperature play important roles in thermodynamics and statistical physics. In fact, using the constant $T_{\rm{KH}}$ model, we can examine relaxation processes for a universe at constant temperature on a dynamic horizon, as if the dynamic horizon is in contact with a heat bath. In this sense, the constant $T_{\rm{KH}}$ model should extend the concept of horizons at constant temperature. The constant $T_{\rm{KH}}$ model should be a good model for studying the relaxation processes for the universe at constant $T_{\rm{KH}}$ Koma19 .
@@
In this section, we observed typical evolutions of the thermodynamic quantities in the modified thermodynamic relation using the $\Lambda$CDM model and the constant $T_{\rm{KH}}$ model. The constant $T_{\rm{KH}}$ model is simply one viable scenario with a constant horizon temperature Koma19 . However, the results for the constant $T_{\rm{KH}}$ model will contribute to the study of thermodynamics and statistical physics on dynamic horizons because the horizon temperature is constant. For example, a constant $T_{\rm{KH}}$ universe always satisfies the holographic equipartition law of energy. Therefore, the emergence of cosmic space based on the law Padma2010 ; Verlinde1 ; HDE ; Padmanabhan2004 ; ShuGong2011 ; Koma14 ; Koma15 ; Koma16 ; Koma17 ; Koma19 ; Koma20 ; Padma2012AB ; Cai2012 ; Hashemi ; Moradpour ; Wang ; Koma10 ; Koma11 ; Koma12 ; Koma18 ; Krishna20172019 ; Mathew2022 ; Chen2022 ; Luciano ; Mathew2023 ; Mathew2023b ; Pad2017 ; Tu2018 ; Tu2019 ; Chen2024 , which can lead to cosmological equations, should be discussed from a different viewpoint, using the modified thermodynamic relation. Of course, in general, the modified thermodynamic relation is given by Eq. (29) and the cosmological equations are given by Eqs. (1)–(3). In this case, we should examine the cosmological constant problem, by applying the second law of thermodynamics to the equations and extending a method used in previous works Koma11 ; Koma12 . These tasks are left for future research.
The modified thermodynamic relation examined in this paper should help to understand the properties of various cosmological models from a thermodynamic viewpoint. The present study should provide new insights into the discussion of thermodynamic cosmological scenarios.